# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v2

@@ IV-C RQ2: Efficiency in Computational Cost
Table V: Comparison of the average generation time per problem (in seconds) by Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP (the brackets indicate the change ratio relative to Greedy)
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 1.18 2.08 (+76.79%) 1.51 (+28.33%) 1.84 (+56.35%) 1.14 1.43 (+25.97%) 1.15 (+1.54%) 1.77 (+55.76%)
DS-6.7B 2.47 6.27 (+154.15%) 3.54 (+43.53%) 4.10 (+66.40%) 2.12 2.97 (+40.03%) 2.28 (+7.62%) 2.64 (+24.38%)
ST-3B 1.50 2.87 (+91.14%) 1.95 (+29.73%) 2.18 (+45.27%) 1.17 1.48 (+26.57%) 0.99 (-15.53%) 1.35 (+15.83%)
CL-7B 5.02 13.69 (+172.58%) 5.72 (+13.85%) 6.41 (+27.75%) 3.23 5.28 (+63.21%) 2.65 (-18.10%) 3.65 (+12.93%)
QW-0.6B 3.67 3.04 (-16.93%) 3.50 (-4.41%) 3.17 (-13.45%) 6.92 2.07 (-70.10%) 4.76 (-31.12%) 5.72 (-17.35%)
QW-1.7B 2.70 2.71 (+0.60%) 3.34 (+23.71%) 4.05 (+50.31%) 2.83 2.57 (-9.30%) 2.84 (+0.16%) 5.46 (+92.86%)
QW-4B 4.19 7.20 (+72.05%) 7.32 (+74.89%) 7.44 (+77.66%) 2.71 2.10 (-22.62%) 4.03 (+48.81%) 5.15 (+89.99%)
QW-8B 5.05 9.37 (+85.51%) 7.01 (+38.77%) 10.12 (+100.46%) 4.03 3.44 (-14.74%) 4.85 (+20.23%) 5.69 (+41.00%)
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: AdaDec selectively pauses decoding at a low rate (3.95–11.88%), reducing token count by up to 54.72% and frequently lowering latency, thereby demonstrating both efficient and accurate decoding.
IV-D RQ3: Ablation Study
We perform an ablation study to evaluate the key components of AdaDec. Specifically, we first investigate the effectiveness of the learned entropy threshold $\tau^{\bm{LM}}$ in identifying top-1 ground-truth tokens based on entropy. Then, we compare two variants of AdaDec to analyze the impact of different lookahead length selection strategies.
@@ IV-D RQ3: Ablation Study
Effectiveness of Learned Entropy Thresholds. Table VI summarizes the performance of a logistic regression classifier trained with learned entropy thresholds ($\tau^{LM}$) across various language models. Overall, the classifier demonstrates strong effectiveness in predicting whether the top-1 token matches the ground truth. The highest performance is observed on DS-6.7B, achieving an accuracy of 91.05%, an F1 score of 0.9512, and an AUC of 0.9305. Other models, including DS-1.3B, ST-3B, and CL-7B, also perform well, with F1 scores above 0.94 and AUC values exceeding 0.89. Notably, CL-7B exhibits the highest learned entropy threshold ($\tau^{LM}=1.9353$), suggesting a greater tolerance. These results support the effectiveness of learned entropy thresholds as a reliable and lightweight signal for triggering pauses and reranking.
@@
Table VI: Evaluation of logistic regression models trained to predict whether the ground truth token is the top-1 prediction. The learned $\bm{\tau}^{\bm{LM}}$ denotes the entropy threshold selected via logistic regression.
@@
Impact of Lookahead Length Selection Strategies. Both $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ consistently improve pass@1 accuracy over Greedy decoding across all tested models on HumanEval and MBPP benchmarks. On HumanEval, $\textsc{AdaDec}_{L=5}$ achieves a slightly higher average improvement (+4.42%) compared to $\textsc{AdaDec}_{dynL}$ (+3.96%), with $\textsc{AdaDec}_{L=5}$ outperforming $\textsc{AdaDec}_{dynL}$ on four out of eight models and achieving parity on one. Conversely, on MBPP, $\textsc{AdaDec}_{dynL}$ attains the highest average gain (+6.12%) versus $\textsc{AdaDec}_{L=5}$ (+5.69%), outperforming $\textsc{AdaDec}_{L=5}$ on four models and matching it on three. Looking at individual models, $\textsc{AdaDec}_{L=5}$ excels on DS-1.3B and QW-8B on HumanEval, showing the largest relative gains, while $\textsc{AdaDec}_{dynL}$ performs better on QW-4B and CL-7B. On MBPP, $\textsc{AdaDec}_{dynL}$ slightly edges out $\textsc{AdaDec}_{L=5}$ on larger models such as DS-6.7B and QW-8B.
Overall, $\textsc{AdaDec}_{L=5}$ tends to deliver stronger improvements on HumanEval, particularly for smaller to mid-sized models, whereas $\textsc{AdaDec}_{dynL}$ shows more consistent gains on MBPP, especially with larger models. This suggests that $\textsc{AdaDec}_{L=5}$’s fixed lookahead length may better suit certain tasks or models, while $\textsc{AdaDec}_{dynL}$’s adaptive lookahead length offers advantages on others, reflecting a tradeoff between stability and flexibility in lookahead strategy.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: The learned entropy threshold $\tau^{\bm{LM}}$ effectively distinguishes uncertain decoding steps, enabling accurate, adaptive lookahead decisions. Both $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ improve pass@1 over Greedy decoding, with $\textsc{AdaDec}_{L=5}$ performing better on HumanEval and $\textsc{AdaDec}_{dynL}$ showing stronger gains on MBPP, reflecting a tradeoff between fixed and dynamic lookahead strategies.
IV-E RQ4: Generalizability across Models and Datasets
We evaluate the generalizability of AdaDec along two dimensions:
Cross-Model Generalizability. We evaluate AdaDec across a broad spectrum of open-source LLMs ranging from 0.6B to 8B parameters, including code-specialized models like DeepSeek-Coder (1.3B, 6.7B), CodeLlama-7B, and StableCode-3B, as well as general-purpose models such as Qwen3 (0.6B, 1.7B, 4B, 8B). In all cases, AdaDec consistently improves Pass@1 accuracy, demonstrating strong robustness across architectures and scales.
Cross-Dataset Generalizability. We further assess AdaDec on two diverse code generation benchmarks, HumanEval and MBPP, which differ in problem complexity and code structure. As shown in Table II, AdaDec increases Pass@1 by an average of around 4% on HumanEval and around 6% on MBPP. These results confirm AdaDec’s effectiveness across a range of programming tasks.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: AdaDec generalizes well across models and datasets, consistently improving Pass@1 accuracy on LLMs from 0.6B to 8B and achieving reliable gains on both HumanEval and MBPP.
V Discussion
V-A Limitations
@@
Limitations of $\textsc{AdaDec}_{dynL}$. The $\textsc{AdaDec}_{dynL}$ variant leverages explicit line breaks in Python code to determine dynamic lookahead length. While effective in this context, such structural cues may not generalize well to other domains—such as natural language generation—or to programming languages that lack clear or consistent line delimiters.
@@
V-B Case Study
To demonstrate how AdaDec enhances decoding decisions in practice, we present a representative example from the HumanEval benchmark (Figure 5). The task is to implement the function check_dict_case, which returns True if all dictionary keys are either lowercase or uppercase, and False otherwise. The prompt provides detailed instructions along with multiple illustrative examples.
Figure 5: A case study from HumanEval
@@ V-B Case Study
VI Threats to Validity
Internal Validity. A key threat to internal validity lies in identifying drift points in our empirical study. We use a semi-automated AST structural comparison pipeline followed by human verification. While this reduces manual annotation effort, it may still miss logic-altering changes caused by identifier drift, leading to false negatives. To address this, we further conduct RQ2 in our empirical study to validate the correlation between entropy and ground-truth token ranks.
External Validity. A major threat to external validity is dataset diversity. We evaluate AdaDec on two widely used Python benchmarks—HumanEval and MBPP—which focus on function synthesis. The generalizability of our approach to other languages or larger, more complex codebases remains untested. Another potential threat is the exclusion of closed-source or commercial models. While we evaluate several open-source models (DeepSeek, StableCode, CodeLlama, Qwen3), our findings may not fully extend to proprietary models such as GPT-4 or Codex.
VII Related Work
VII-A LLM-based Code Generation
@@
Large Language Models (LLMs) have shown strong potential in software engineering tasks such as code generation [1, 2, 3, 4, 5], unit test generation [27, 28, 29, 30, 31], and comment generation [32, 33, 34]. Typically, LLMs generate code from textual requirements via prompt-based interaction. By learning expressive representations from large-scale code and documentation corpora, they can produce syntactically correct and context-aware code snippets [8, 9, 10]. Open-foundation models further advance tasks like code completion, bug fixing [11, 12, 13], and test-case generation.
@@
To evaluate their effectiveness, benchmarks like HumanEval [16], MBPP [26], and DS-1000 [35] focus on function-level generation, while ClassEval [36] extends evaluation to class-level tasks. More recently, CoderEval [37] and DevEval [38] target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.
VII-B Decoding strategies for LLMs
@@
Large Language Models (LLMs) generate text in an autoregressive manner, producing one token at a time. At each generation step, the model samples a token from its predicted probability distribution. The most straightforward sampling strategy is greedy decoding, which selects the token with the highest probability as the next word. A more refined variant is beam search [39], which maintains multiple candidate sequences at each decoding step and preserves the top-k sequences based on cumulative log-probabilities.
@@
Beyond these, commonly used sampling strategies include temperature sampling [40], top-p (nucleus) sampling [40], and top-k sampling [40], each designed to balance diversity and coherence in generated outputs.
@@
To further improve sampling for code generation tasks, Li et al. introduced AdapT Sampling [41] (Adaptive Temperature Sampling). This method dynamically adjusts the temperature parameter T based on token context. For instance, when the current token initiates a code block, a higher T is used to encourage diversity and creativity. In contrast, for other tokens, a lower T is applied to suppress random noise and preserve syntactic and semantic correctness.
@@
Another relevant approach is UnCert-CoT [42], which enhances code generation through uncertainty-aware Chain-of-Thought (CoT) reasoning, activating multi-path reasoning when token uncertainty surpasses a fixed threshold. While effective, its relies on fixed thresholds and costly CoT reasoning. In contrast, AdaDec learns model-specific uncertainty thresholds for stronger generalization, while its lightweight rerank mechanism cuts computational costs significantly without compromising benchmark performance.
VIII Conclusion
@@
This paper presents AdaDec, an uncertainty-guided adaptive decoding framework for large language model-based code generation. Motivated by an empirical study revealing that ranking errors at high-uncertainty decoding steps often lead to logic drift, AdaDec introduces a novel pause-then-rerank mechanism that dynamically identifies and corrects uncertain token predictions using Shannon entropy. By learning model-specific entropy thresholds and employing a lookahead-based reranking strategy, AdaDec achieves substantial improvements in both accuracy and efficiency. Experiments across HumanEval and MBPP benchmarks demonstrate that AdaDec consistently outperforms greedy decoding and competes favorably with beam search while significantly reducing decoding overhead. These findings underscore the importance of token-level uncertainty modeling and open up new directions for adaptive decoding in code generation and beyond.

### v3

@@ 4.3. RQ2: Efficiency in Computational Cost
Table 4. Comparison of the average generation time per problem (in seconds) by Greedy, AdapT, AdaDec on HumanEval+, MBPP+, DevEval (the brackets indicate the change ratio relative to Greedy)
Model HumanEval+ MBPP+ DevEval
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 1.15 2.42 1.46 0.54 0.63 0.97 4.74 9.61 10.24
DS-6.7B 2.37 3.68 3.44 1.02 4.24 1.30 15.54 11.55 26.08
ST-3B 1.70 3.02 2.06 0.59 0.81 0.76 6.87 8.00 16.72
QW-0.6B 6.22 4.07 5.25 10.02 6.14 2.89 20.02 13.86 21.18
QW-1.7B 8.62 4.23 5.27 2.39 2.54 2.28 18.21 13.96 30.33
QW-4B 5.41 7.57 8.06 4.09 3.76 3.37 28.17 17.94 43.97
QW-8B 5.39 4.73 6.94 2.84 4.83 4.01 64.56 16.80 68.08
AVG. $\Delta$ -0.16 +0.23 +0.21 -0.84 -9.48 +8.36
Summary: AdaDec maintains high efficiency by selectively applying pause-and-rerank at a small fraction of high-uncertainty steps, with pause rates mostly below 10%. While this targeted lookahead adds moderate latency compared to Greedy, it remains manageable and effectively balances quality and efficiency. AdapT can sometimes reduce runtime via stochastic sampling, but AdaDec delivers more consistent improvements in functional correctness.
4.4. RQ3: Ablation Study
We perform an ablation study to evaluate the key components of AdaDec. In particular, we investigate whether the learned entropy threshold $\tau^{\bm{LM}}$ serves as an effective signal, both in terms of predictive ability (as a classifier) and in practical decoding performance.
@@ 4.4. RQ3: Ablation Study
Predictive Effectiveness of Learned Entropy Thresholds. Table 5 summarizes the performance of a logistic regression classifier trained with learned entropy thresholds ($\tau^{\bm{LM}}$) across various language models. Overall, the classifier demonstrates strong effectiveness in predicting whether the top-1 token matches the ground truth. The highest performance is observed on DS-6.7B, achieving an accuracy of 91.05%, an F1 score of 0.9512, and an AUC of 0.9305. Other models, including DS-1.3B and ST-3B, also perform well, with F1 scores above 0.94 and AUC values exceeding 0.89. These results support the effectiveness of learned entropy thresholds as a reliable and lightweight signal for triggering pauses and reranking.
@@
Table 5. Evaluation of logistic regression models trained to predict whether the ground truth token is the top-1 prediction. The learned $\bm{\tau}^{\bm{LM}}$ denotes the entropy threshold selected via logistic regression.
@@
AVG. $\Delta$ 0.8852 0.9117 0.9591 0.9348 0.9120 0.9393
Practical Effectiveness of Learned Entropy Thresholds. To further assess the impact of learned thresholds in practice, we compare the decoding performance of models using the learned threshold versus a fixed threshold. The fixed threshold is set to 1.2. Table 6 reports pass@1 results on HumanEval+. The experiments show that using the learned thresholds consistently improves performance compared to the fixed threshold. For DS-1.3B, DS-6.7B, ST-3B, and QW-0.6B, the learned thresholds are close to 1.2, resulting in modest gains. In contrast, for QW-1.7B, QW-4B, and QW-8B, the learned thresholds differ more from 1.2, yielding more noticeable improvements in decoding performance. This confirms that the learned thresholds not only perform well in predictive tasks but also translate into tangible gains during actual decoding.
Table 6. Comparison of Learned vs. Fixed Entropy Thresholds on HumanEval+ (pass@1). Fixed threshold is set to 1.2.
Strategy DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B AVG
Fixed 57.32% 65.24% 50.00% 21.34% 41.46% 51.22% 61.59% 49.74%
Learned 57.32% 65.24% 50.61% 22.56% 44.51% 56.10% 62.80% 51.31%
$\Delta$ 0.00% 0.00% +0.61% +1.22% +3.05% +4.88% +1.21% +1.57%
Summary: The learned entropy threshold $\tau^{\bm{LM}}$ is both theoretically sound and practically effective: it reliably distinguishes uncertain decoding steps and improves decoding performance when compared to fixed-threshold baselines.
4.5. RQ4: Parameter Sensitivity
To investigate the impact of lookahead length $L$ on decoding performance, we conduct experiments on HumanEval+ using DS-1.3B as the backbone model. We vary $L$ from 2 to 9 and report pass@1 results. Figure 6 illustrates the trend.
Figure 6. Pass@1 Performance of the DS-1.3B Model on the HumanEval+ Dataset Under Different Lookahead Length (L) Configurations.
The results show that pass@1 increases steadily with larger $L$ up to $L=5$, at which point it reaches the peak performance. Beyond $L=5$, the performance declines slightly and then plateaus. This behavior can be explained by the trade-off between local guidance and global noise: shorter lookaheads may fail to capture sufficient future context, while excessively long lookaheads introduce noise and dilute the relevance of local signals. Importantly, across all values of $L$, AdaDec consistently outperforms the baseline decoding strategies, demonstrating the robustness of our approach to the choice of lookahead length.
Summary: The lookahead length $L$ exhibits a non-monotonic effect on performance: moderate values (e.g., $L=5$) maximize the benefits of lookahead guidance, while excessively small or large values reduce effectiveness. Notably, AdaDec consistently surpasses baseline performance regardless of the choice of $L$.
5. Discussion
5.1. Limitations
@@
Potential Bias from Training Data. Although we carefully avoided direct overlap between training and evaluation datasets, our threshold learning still relies on synthetic data generated from a specific corpus. If the distribution of training data diverges significantly from the deployment scenario, the learned thresholds may fail to transfer robustly. Future work could explore continual adaptation or online threshold adjustment.
@@
5.2. Case Study
To demonstrate how AdaDec enhances decoding decisions in practice, we present a representative example from the HumanEval+ benchmark (Figure 7). The task is to implement the function check_dict_case, which returns True if all dictionary keys are either lowercase or uppercase, and False otherwise. The prompt provides detailed instructions along with multiple illustrative examples.
Figure 7. A case study from HumanEval+
@@ 5.2. Case Study
6. Threats to Validity
We analyze threats to validity across four dimensions—internal, external, construct, and conclusion—and note our mitigation strategies or key limitations.
Internal Validity. Drift point identification may be imperfect: our semi-automated AST comparison could miss subtle semantic shifts (e.g., identifier drift), and human verification introduces potential annotator bias. We mitigated these risks through clear guidelines and cross-checking. Entropy-based analysis may also depend on tokenizer or model-specific distributions; to reduce this risk, we validated results across multiple models and confirmed entropy–rank correlations in RQ2.
External Validity. We focus on open-source models (DeepSeek, StableCode, Qwen3) and three benchmarks (HumanEval+, MBPP+, DevEval). Results may not fully generalize to proprietary models (e.g., GPT-4, Codex) or to broader real-world tasks. Future work should extend evaluation to more diverse datasets and model families.
Construct Validity. We evaluate functional correctness (Pass@1) and efficiency (latency), which directly measure the primary objectives of correctness and performance. Nevertheless, these metrics do not capture other important qualities such as readability or adherence to coding conventions. Similarly, our use of Shannon entropy as an uncertainty signal assumes it faithfully reflects model uncertainty, though it may only provide a coarse approximation. Alternative measures (e.g., calibration error, variance across sampling runs) could offer complementary insights.
Conclusion Validity. Conclusion validity threats arise from the statistical robustness of our empirical results. Although we report consistent trends across multiple models and datasets, the sample size of benchmarks such as HumanEval+ remains relatively small, which may limit the statistical power of our findings. Additionally, hyperparameter choices (e.g., lookahead length $L$, candidate pool size $B$) may affect performance outcomes, and while we explored reasonable ranges, further sensitivity analysis would strengthen confidence in our conclusions.
7. Related Work
7.1. LLM-based Code Generation
@@
Large Language Models (LLMs) have shown strong potential in software engineering tasks such as code generation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024), unit test generation (Nan et al., 2025; Yang et al., 2024; Yuan et al., 2024; Lops et al., 2025; Ryan et al., 2024), and comment generation (Haider et al., 2024; Katzy et al., 2025; Lu and Liu, 2024). Typically, LLMs generate code from textual requirements via prompt-based interaction. By learning expressive representations from large-scale code and documentation corpora, they can produce syntactically correct and context-aware code snippets (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). Open-foundation models further advance tasks like code completion, bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation.
@@
To evaluate their effectiveness, benchmarks like HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), and DS-1000 (Lai et al., 2022) focus on function-level generation, while ClassEval (Du et al., 2023) extends evaluation to class-level tasks. More recently, CoderEval (Zhang et al., 2024) and DevEval (Li et al., 2024) target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.
7.2. Decoding strategies for LLMs
@@
Large Language Models (LLMs) generate text in an autoregressive manner, producing one token at a time. At each generation step, the model samples a token from its predicted probability distribution. The most straightforward sampling strategy is greedy decoding, which selects the token with the highest probability as the next word. A more refined variant is beam search (Freitag and Al-Onaizan, 2017), which maintains multiple candidate sequences at each decoding step and preserves the top-k sequences based on cumulative log-probabilities.
@@
Beyond these, commonly used sampling strategies include temperature sampling (Holtzman et al., 2020), top-p (nucleus) sampling (Holtzman et al., 2020), and top-k sampling (Holtzman et al., 2020), each designed to balance diversity and coherence in generated outputs.
@@
To further improve sampling for code generation tasks, Li et al. introduced AdapT Sampling (Zhu et al., 2023) (Adaptive Temperature Sampling). This method dynamically adjusts the temperature parameter T based on token context. For instance, when the current token initiates a code block, a higher T is used to encourage diversity and creativity. In contrast, for other tokens, a lower T is applied to suppress random noise and preserve syntactic and semantic correctness.
@@
Another relevant approach is UnCert-CoT (Zhu et al., 2025), which enhances code generation through uncertainty-aware Chain-of-Thought (CoT) reasoning, activating multi-path reasoning when token uncertainty surpasses a fixed threshold. While effective, its relies on fixed thresholds and costly CoT reasoning. In contrast, AdaDec learns model-specific uncertainty thresholds for stronger generalization, while its lightweight rerank mechanism cuts computational costs significantly without compromising benchmark performance.
8. Conclusion
@@
This paper presents AdaDec, an uncertainty-guided adaptive decoding framework for large language model-based code generation. Motivated by an empirical study revealing that ranking errors at high-uncertainty decoding steps often lead to logic drift, AdaDec introduces a novel pause-then-rerank mechanism that dynamically identifies and corrects uncertain token predictions using Shannon entropy. By learning model-specific entropy thresholds and employing a lookahead-based reranking strategy, AdaDec achieves substantial improvements in accuracy while maintaining computational cost and efficiency within acceptable bounds. Experiments on HumanEval+, MBPP+, and DevEval benchmarks demonstrate that AdaDec consistently outperforms all compared methods. These findings underscore the importance of token-level uncertainty modeling and open up new directions for adaptive decoding in code generation and beyond.
9. Data Availability
To facilitate the replication study, we have released our data and code at : https://anonymous.4open.science/r/AdaDec-657F.